# CoDiff: Conditional Diffusion Model for Collaborative 3D Object Detection

### v2

@@ Abstract
Collaborative 3D object detection holds significant importance in the field of autonomous driving, as it greatly enhances the perception capabilities of each individual agent by facilitating information exchange among multiple agents. However, in practice, due to pose estimation errors and time delays, the fusion of information across agents often results in feature representations with spatial and temporal noise, leading to detection errors. Diffusion models naturally have the ability to denoise noisy samples to the ideal data, which motivates us to explore the use of diffusion models to address the noise problem between multi-agent systems. In this work, we propose CoDiff, a novel robust collaborative perception framework that leverages the potential of diffusion models to generate more comprehensive and clearer feature representations. To the best of our knowledge, this is the first work to apply diffusion models to multi-agent collaborative perception. Specifically, we project high-dimensional feature map into the latent space of a powerful pre-trained autoencoder. Within this space, individual agent information serves as a condition to guide the diffusion model’s sampling. This process denoises coarse feature maps and progressively refines the fused features. Experimental study on both simulated and real-world datasets demonstrates that the proposed framework CoDiff consistently outperforms existing relevant methods in terms of the collaborative object detection performance, and exhibits highly desired robustness when the pose and delay information of agents is with high-level noise. The code is released at https://github.com/HuangZhe885/CoDiff.
@@ I INTRODUCTION
3D object detection [1, 2, 3, 4, 5, 6] is a fundamental task in autonomous driving, primarily focusing on the localization and identification of specific vehicles in real-world scenarios [7]. It is also attracting considerable interest in domains such as as drones, robots, and the metaverse [8, 9, 10, 11, 12]. However, 3D object detection with a single agent has inherent limitations, such as occlusion and distant objects [13, 6, 14]. Recently, researchers have addressed these limitations by enabling multiple agents to share complementary perception information through communication, leading to more comprehensive and accurate detection. Cooperative 3D object detection [15, 16, 17, 18] is an important field that has made significant progress in terms of high-quality datasets and methods. However, it also faces many challenges, including time delays [19], pose misalignment [17], communication bandwidth limitations [16], and adversarial attacks [20].
To effectively share information, multiple agents need to transmit real-time, precisely aligned features to synchronize their data within a consistent timestamp and spatial coordinate system, which is the foundation for maintaining effective collaboration. However, in real-world scenarios, this information may suffer from time delays and spatial misalignment due to the following reasons: 1) Unstable communication between agents, such as congestion and interruptions, leading to time asynchrony issues; 2) The 6-degree-of-freedom (DoF) poses estimated by each agent’s localization module are imperfect, resulting in unavoidable relative pose errors. These factors severely affect the reliability and quality of information exchange between agents, leading to inaccurate features during fusion. Some previous studies have considered various methods to address these issues. CoAlign [21] introduces a novel agent-object pose graph modeling approach to enhance pose consistency between collaborative agents. RoCo [17] designs an object matching and optimization strategy to correct pose errors. V2XViT [22] incorporates latency as an input for feature compensation, while CoBEVFlow [19] introduces the concept of feature flow to handle irregular time delays. However, in practice, time delays and pose errors may usually coexist and entangle with each other, and existing methods are unable to address both issues simultaneously. As a result, they still produce feature representations with harmful noise, hindering current collaborative perception systems from reaching their full potential.
Figure 1: Illustration of the proposed robust collaborative perception system with the conditional diffusion model, where $q$ is the diffusion process and $p_{\theta}$ is the reverse process.
To address these limitations, we propose a novel hybrid collaborative framework called CoDiff, which leverages Conditional Diffusion Probabilistic Models (CDPM) to address spatial and temporal noise issues in information sharing, thereby improving the quality of multi-agent fused features. In CoDiff, we treat the time delays and pose errors between multiple agents as a unified noise to be learned and handled. We employed diffusion models for feature fusion, replacing the existing feature fusion methods based on regression models and attention mechanisms. CoDiff is primarily consists of two components: perception compression and conditional generation. The perception compression module compresses features into a latent space which makes the computation more efficient while well retaining the perception information. High-resolution feature generation is then performed within this space, well reducing the complexity of both training and sampling processes . The conditional generation module takes the features from different collaborating agents as input and incrementally refine the noisy features, as illustrated in Fig. 1. CoDiff has three main advantages: i) It can adaptively establish relationships between single-agent features and the final fused features, contributing a more compact perception representation; ii) It is capable of learning the distribution resulted from the mixture of different types of noise, mitigating the impact of time delays and pose errors on collaborative information; iii) Unlike existing methods that design specialized fusion modules, our generative approach naturally prevents the introduction of additional noise into the features.
@@
We conduct extensive experiments on both simulation and real-world datasets, including V2XSet[22], OPV2V [23] and DAIR-V2X[24]. Results show that CoDiff consistently achieves the best performance in the task of collaborative 3D object detection with the presence of pose errors and time delays. In summary, the main contributions of this work are:
@@
  • We propose CoDiff, a novel LiDAR-based robust multi-agent collaborative 3D object detection framework that addresses the problem of noisy feature map fusion between multiple agents. To the best of our knowledge, CoDiff is the first work to use a diffusion model to solve time delays and pose inaccuracies in collaborative perception.
  • The proposed CoDiff conducts perception compression to obtain a low-dimensional and computationally more efficient latent space. Within this space, CoDiff generates and refines multi-agent features, ultimately producing a globally fused feature map.
  • Extensive experiments have shown that CoDiff achieves more accurate and robust 3D object detection performance even in scenarios with significant pose noise and large time delays.
@@ II Related Work
Diffusion Models for Perception Tasks. Diffusion models [10, 11, 12, 25], known for their powerful denoising capabilities, have attracted considerable interest in domains such as natural language processing, text transformation, and multimodal data production. Recently, an increasing number of researchers have started investigating the use of diffusion models in perception tasks. DDPM-Segmentation [26] is the initial study that uses diffusion models for semantic segmentation. On the other hand, DiffusionDet [27] considers object detection as a noise-to-box task, aiming to provide accurate object bounding boxes by progressively denoising randomly generated proposals. Diffusion-SS3D [28] utilises the powerful properties of diffusion models in a semi-supervised 3D object detection framework, with the goal of generating more dependable pseudo-labels. Expanding upon this, Diff3DETR [29] extends the approach of Diffusion-SS3D by proposing the first diffusion-based DETR framework. In addition, DifFUSER [30] utilises the noise reduction capabilities of diffusion models to address noise for multimodal fusion (such as image, text, video). In this work, we are motivated to further explore the potential of employing the diffusion model to generate a high-quality representation, proposing the first diffusion-based multi-agent cooperative perception 3D detection method.
@@
Collaborative 3D Object Detection. Collaborative 3D object detection [31, 21, 32, 22, 23] is a particular application in multi-agent systems that allows multiple agents to share information to overcome the inherent limitations of single-agent perception. Many researchers have designed various methods to enhance the perception performance and robustness of collaborative perception systems. In terms of perception performance, the study in the literature [32, 22, 23, 16] implemented a transformer architecture to aggregate information from different agents. In terms of robustness, CoBEVFlow[19] creates a synchrony-robust collaborative system that aligns asynchronous collaboration messages sent by various agents using motion compensation. HEAL[15] smoothly integrates emerging heterogeneous agent types into collaborative perception tasks. CoAlign [21] uses an agent-object pose graph to address pose inaccuracies, and RoCo [17] tackles pose errors with object matching and graph optimization techniques. However, we observe that these methods tend to address only a single aspect of the problem and fail to ensure both perceptual performance and robustness of the model under multiple noise conditions. In contrast, our CoDiff can simultaneously address the effects of various types of noise, such as time delays and pose errors, thereby further improving the accuracy of collaborative 3D object detection.
Figure 2: Overview of the proposed CoDiff framework. Single-agent features are transmitted to the ego vehicle. These features are concatenated and sent into the compression model $\mathcal{E}$, before entering the latent diffusion model. The diffusion model generates denoised features in latent space, which are then decompressed to obtain the denoised multi-agent fused features.
III Problem Definition and Preliminaries
III-A Collaborative 3D object detection.
Assuming there are $N$ agents in the scene, given the point clouds of these agents as input, the goal of collaborative 3D object detection is to classify and locate the 3D bounding boxes of objects within the scene through the cooperation of the agents. This process can be formally expressed as:
@@ III Problem Definition and Preliminaries
$F^{\prime}_{i}$ $=\Phi_{Agg}\left(F_{i},\{\mathcal{M}_{j\rightarrow i}\}_{j=1,2,...,N;j\neq{i}}\right),$ (1b)
$\mathcal{O}_{i}$ $=\Phi_{Dec}\left(F^{\prime}_{i}\right).$ (1c)
Within the collaborative perception framework, each agent $i$ can extract features $F_{i}$ from the raw point cloud observations $\mathcal{X}_{i}$ using an encoding network as in Step (1a). $\mathcal{M}_{j\rightarrow i}$ in Step (1b) denotes the collaboration message sent from agent $j$ to agent $i$. After receiving all the $(N-1)$ messages, agent $i$ will aggregate its feature $F_{i}$ with these messages via a fusion network, producing a fused feature $F^{\prime}_{i}$ as in Step (1b). Finally, Step (1c) uses a decoding network to convert $F^{\prime}_{i}$ to the final perception output $\mathcal{O}_{i}$. As seen above, provided that all networks are well trained, the fused features $F^{\prime}_{i}$ in Step (1b) will heavily rely on the quality of the information $\mathcal{M}_{j\rightarrow i}$ shared between the agents. However, in the real world, time asynchrony and pose estimation errors between agents are inevitable, leading to issues such as time delays and inaccurate pose estimates. These problems make it difficult to achieve accurate feature fusion. In this paper, we propose a novel approach that utilizes a generative model $\Phi_{diff}$ to replace step 1(b) by generating more fine-grained and detailed features $F_{i}^{\prime}=\Phi_{diff}\left(F_{i},\left\{M_{j\rightarrow i}\right\}_{j=1,2,...,N;j\neq{i}}\right)$ , thereby achieving more precise perception results.
III-B Diffusion Model.
Diffusion models are generative models inspired by nonequilibrium thermodynamics [33, 34], designed to learn a data distribution $P\left(x\right)$ by gradually denoising a normally distributed variable. The diffusion forward process transforms an initial data sample $x_{0}$ into a noisy sample $x_{t}$ at time $t$ by introducing noise that is determined by the noise variance schedule $\beta_{t}$. For convenience, we denote a series of constant: $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=1}^{t}a_{s}$. The forward process to generate a noisy sample $x_{t}$ from $x_{0}$ can be defined by
where $\mathcal{N}$ denotes a Gaussian distribution and $\mathbf{I}$ is an identity matrix. A neural network is then trained to reverse this diffusion process. This network can be interpreted as an equally weighted sequence of denoising autoencoders $\epsilon_{\theta}\left(x_{t},t\right);t=1,...,T$, which are trained to predict a denoised variant of their input $x_{t}$, where $x_{t}$ is a noisy version of the input $x$. The corresponding objective can be simplified to
Diffusion models can learn diverse data distribution in multiple domains. We aim to leverage this advantage and design more suitable generative conditions to address the issues of time delays and pose inaccuracies between agents in collaborative perception.

### v3

# CoDiff: Robust Collaborative 3D object deteciton via Diffusion-Based Feature Realignment

@@ Abstract
Collaborative 3D object detection holds significant importance in the field of autonomous driving, as it greatly enhances the perception capabilities of each individual agent by facilitating information exchange among multiple agents. Most collaborative perception systems rely on an idealized noise-free assumption, ignoring delays and pose inconsistencies during interaction. However, real-world deployment inevitably introduces noise, causing severe feature misalignments and performance degradation. Diffusion models naturally have the ability to denoise noisy samples to the ideal data, which motivates us to explore the use of diffusion models to address feature misalignment in multiagent systems caused by noise. In this work, we propose CoDiff, a novel robust collaborative perception framework that leverages the potential of diffusion models to generate more comprehensive and clearer feature representations. Specifically, CoDiff consists of two core modules: a Pose Calibration Module (PCM), which establishes associations between objects detected by different agents and optimizes their relative poses by minimizing alignment errors; and a Time Calibration Module (TCM), which leverages single-agent features as conditions to remove delay-induced noise and generate clearer and more comprehensive collaborative representations. Experimental study on both simulated and real-world datasets demonstrates that the proposed framework CoDiff consistently outperforms existing relevant methods in terms of the collaborative object detection performance and exhibits highly desired robustness when the pose and delay information of agents is with high-level noise. Ablation studies further validate the effectiveness of key components and design choices within the framework.
Index Terms:
Collaborative Perception, 3D Point cloud, Graph Matching, Diffusion Model, 3D object detection
@@ I Introduction
3D object detection [1, 2, 3, 4, 5, 6] is a fundamental task in autonomous driving. It has gained significant interest in domains including drones, robotics, and the metaverse [7, 8, 9, 10, 11, 12]. However, 3D object detection with a single agent has inherent limitations, such as occlusion and missing detection of distant objects [13, 14, 15, 16]. Recently, researchers have addressed these limitations by enabling multiple agents to share complementary perception information through communication, leading to more comprehensive and accurate detection. Cooperative 3D object detection [17, 18, 19, 20] is an important field that has made significant progress in terms of high-quality datasets and methods. However, it also faces many challenges, including time delays [21], pose inconsistencies [19], communication bandwidth limitations [18], and adversarial attacks [22].
Existing methods have considered various methods to address these challenges. CoAlign [23] introduces a novel agent-object pose graph modeling approach to improve pose consistency between collaborative agents. RoCo [19] designs a matching and optimization strategy to correct pose inconsistencies. V2XViT [24] incorporates latency as an input for feature compensation, while CoBEVFlow [21] introduces the concept of feature flow to handle irregular time delays. The above methods work well when the noise is minimal and the number of vehicles is limited. However, in crowded and complex scenarios, as noise increases, misalignment in object information transmitted by different agents increases, making it increasingly difficult to align features at the same location. This situation ultimately reduces the accuracy of 3D object detection.
Figure 1: Illustration of collaborative perception under noise and the perception results w.o./w. CoDiff. Red boxes are detection results and green boxes are the ground truth.
To effectively share information, multiple agents need to transmit precisely aligned real-time features to synchronize their data within a consistent timestamp and spatial coordinate system, which is the foundation for maintaining effective collaboration. However, in real-world scenarios, we observe that the transmitted information may suffer from pose and time misalignment. As illustrated in Fig.1, two key factors significantly affect the performance of collaborative perception. The first issue is time delays, as shown in Fig.1(a), when agent $j$ sends information to agent $i$, unstable inter-agent communication (e.g. congestion or interruptions) may introduce delays, resulting in misaligned timestamps between agents. Secondly, the 6-DoF pose estimation conducted by each agent’s localization module is inherently imperfect, inevitably leading to relative pose inconsistencies. For instance, in Fig.1(b), a noticeable positional deviation between object 3 detected by agent $i$ (orange box) and object 2 detected by agent $j$ (blue box), which adversely affects subsequent feature fusion. When pose inconsistencies or time delays are large, this misalignment becomes even more pronounced. Therefore, this work focuses on addressing the inter-agent feature misalignment in 3D object detection caused by time delays and pose inconsistencies, with the goal of mitigating the negative impact of such noise on collaborative detection systems.
To address this limitation, we propose a novel hybrid collaborative framework called CoDiff, which can address pose and time noise in information sharing, mitigate feature misalignment, and ultimately enhance detection accuracy. In CoDiff, we treat the time delays between multiple agents as a unified noise to be learned and handled, and employ conditional diffusion models for feature fusion, replacing the existing methods based on regression models and attention mechanisms. CoDiff primarily consists of two components: a Pose Calibration Module (PCM) and a Time Calibration Module (TCM). PCM establishes correspondences among multiple objects and constructs a pose optimization graph to effectively adjust agent pose inconsistencies. TCM first compresses the high-dimensional features of agents into a latent space and then employs a conditional diffusion model to progressively synthesize denoised, higher-quality features, as illustrated in Fig. 2. CoDiff has three main advantages: i) The pose adjustment method in CoDiff is unsupervised, requiring no ground-truth pose annotations for either agents or objects, and it exhibits strong adaptability to various pose inconsistencies; ii) Unlike existing methods that design specialized fusion modules, the proposed fusion approach is a generative method based on diffusion models, which fundamentally prevents the introduction of additional noise into the features; iii) Even in crowded and complex environments, CoDiff can accurately align the agents and improve detection accuracy.
@@
We conducted extensive experiments on both simulation and real-world datasets, including V2XSet[24], OPV2V [25] and DAIR-V2X[26]. Results show that CoDiff consistently achieves the best performance in the task of collaborative 3D object detection with the presence of pose inconsistencies and time delays. In summary, the main contributions of this work are as follows:
@@
  • We propose CoDiff, a novel and powerful diffusion-based collaborative perception framework that addresses pose inconsistencies and time delays among multiple agents, alleviates feature misalignment, and improves collaborative detection performance.
  • CoDiff first establishes object correspondences through graph matching, refines their relative poses by minimizing alignment errors, and subsequently employs a diffusion-based generative fusion strategy to mitigate time delays, ultimately yielding clearer and more comprehensive representations.
  • Extensive experiments have shown that CoDiff consistently enhances the accuracy and robustness of 3D object detection, even under challenging conditions with dense traffic and significant noise.
Figure 2: Illustration of the proposed robust 3D multi-agent collaborative perception system with PCM and TCM module, where $q$ is the diffusion process and $p_{\theta}$ is the reverse process, massage $M_{j\rightarrow i}$ is the condition.
@@ II Related Work
Collaborative 3D Object Detection. Collaborative 3D object detection [27, 23, 28, 24, 25] enables multiple agents to share information to overcome the limitations of single-agent perception, and various methods have been proposed to improve its performance and robustness. In terms of perception performance, the study in the literature [28, 24, 25, 18] implemented a transformer architecture to aggregate information from different agents. To address communication delays, SyncNet[27] utilizes previous multiframe data to augment present information, implementing feature attention co-symbiotic estimation and temporal modulation approaches. In terms of robustness, CoBEVFlow[21] creates a synchrony-robust collaborative system that aligns asynchronous collaboration messages sent by various agents using motion compensation. HEAL[17] smoothly integrates emerging heterogeneous agent types into collaborative perception tasks. CoAlign [23] uses an agent-object pose graph to address pose inaccuracies, and RoCo [19] tackles pose inconsistencies with object matching and graph optimization techniques. However, these methods are generally designed for scenarios with a single type of noise and relatively simple traffic conditions, and their accuracy and robustness are not guaranteed in dense, complex environments. In contrast, CoDiff explicitly accounts for multiple noise sources, including time delays and pose inconsistencies, thereby improving feature fusion quality and ultimately enhancing the accuracy of collaborative 3D object detection.
Diffusion Models for Perception Tasks. Diffusion model (DM) [9, 10, 29, 30], known for their denoising capability, have gained attention in domains such as natural language processing, text transformation, and multimodal data generation, and are increasingly explored in perception tasks. DDPM-Segmentation [31] is the initial study that uses diffusion models for semantic segmentation. On the other hand, DiffusionDet [32] considers object detection as a noise-to-box task, aiming to provide accurate object bounding boxes by denoising randomly generated proposals. Expanding upon this, Diff3DETR [33] extends the approach of Diffusion-SS3D by proposing the first diffusion-based DETR framework. In addition, DifFUSER [34] utilizes the noise reduction capabilities of diffusion models to address noise for multimodal fusion (such as image, text, and video). In order to enable DM training under limited computational resources, Rombach [35] proposed the Latent Diffusion Model (LDM), which can significantly improve training and sampling efficiency without compromising the quality of the denoising diffusion model. In this work, we are motivated to further explore the potential of employing the diffusion model to generate a high-quality representation, proposing a diffusion-based multi-agent cooperative perception 3D detection method.
III Collaborative Perception and Our proposed method
III-A Collaborative Perception
It is assumed that there are $N$ agents (i.e., collaborators) in the scene. For each agent $i$, let $\mathcal{X}_{i}$ denote its input point cloud and $\mathcal{O}_{i}$ denote its corresponding output. Furthermore, $\mathcal{M}_{j\rightarrow i}$ represents the collaboration message transmitted from agent $j$ to agent $i$. The goal of collaborative 3D object detection is to accurately classify and localize the 3D bounding boxes of objects in the scene through the cooperation of multiple agents. This process can be formally defined as follows:
@@ III-A Collaborative Perception
$\mathcal{M}_{j\rightarrow i}$ $=\Phi_{Proj}\left(\xi_{i}^{t},\left(F_{j},\xi_{j}^{t}\right)\right),\quad j=1,\cdots,N;j\neq{i}$ (1b)
$F^{\prime}_{i}$ $=\Phi_{Agg}\left(F_{i},\{\mathcal{M}_{j\rightarrow i}\}_{j=1,2,...,N;j\neq{i}}\right),$ (1c)
$\mathcal{O}_{i}$ $=\Phi_{Dec}\left(F^{\prime}_{i}\right).$ (1d)
Within the collaborative perception framework, each agent $i$ extracts features $F_{i}$ from its raw point cloud observations $\mathcal{X}_{i}$ using an encoding network, as shown in Step (1a). Then, in Step (1b), each agent $j$ employs the projection module $\Phi_{Proj}\left(\cdot\right)$ to transform its feature $F_{j}$ into the coordinate system of agent $i$ based on $\xi_{i}$, $\xi_{j}$, and $t$, where $\xi_{i}$ and $\xi_{j}$ denote the poses of agents $i$ and $j$, and $t$ denotes the timestamp. The projected feature is sent as the collaboration message $\mathcal{M}_{j\rightarrow i}$ to agent $i$. After receiving all (N1)(N-1) messages, agent $i$ aggregates them with its own feature $F_{i}$ using $\Phi{Agg}(\cdot)$ to produce the fused feature $F^{\prime}_{i}$ (Step 1c), which is finally decoded into the perception output $\mathcal{O}_{i}$ (Step 1d).
III-B Issue of feature misalignment
As shown above, assuming all networks are well-trained, the fused features $F^{\prime}_{i}$ in Step (1c) depend heavily on the quality of the information $\mathcal{M}_{j\rightarrow i}$ shared between agents. However, the poses estimated by each agent are not perfect in practice, i.e., $\{\xi_{j,j=1,\cdots,N\}}$ have errors, which adversely affects the accuracy of feature projection. Moreover, time delays may cause inconsistencies in the transmitted feature $\mathcal{M}_{j\rightarrow i}$, leading to feature misalignment when agent $i$ processes the message, which ultimately degrades the performance of 3D object detection.
To address this feature misalignment, existing works have explored different strategies to enhance the quality of fused features across multiple agents, including pose calibration [19], time compensation [21], and feature fusion techniques [24]. However, these methods are designed to handle a single type of error in isolation. In crowded and dynamic environments, agents are simultaneously subject to both pose inaccuracies and communication delays. Under such conditions, this methods are prone to introducing significant misalignment errors during the feature fusion stage. These misalignment errors will in turn degrade localization accuracy and ultimately compromise the overall performance of collaborative perception.
Figure 3: Overview of the CoDiff system. Bounding boxes and poses are transmitted as messages to other agents, where pose calibration is performed through object matching and graph optimization. The calibrated poses are used to transform features into the ego coordinate system, a conditional diffusion model is leveraged to mitigate time delay noise, thereby generating more robust and reliable multi-agent fused features.